If 4 apples and 2 oranges equals $1 and 2 apples and 3 oranges equals $070, how
much does each apple and each orange cost?

Answer by mananth(16949)

let each apple cost x
each orange cost y
..
4x+2y=1...................1
2x+3y = 0.70..............2
..
multiply (2) by -2
-4x-6y=-1.40
add this to (1)
4x-4x+2y-6y=1-1.40
-4y=-0.4
/-4
y=-0.4/-4
y=0.1 = 10 cents cost of orange
plug value of y in (1)
4x+2y=1
4x+2*0.1=1
4x+0.2=1
-0.2
4x=1-0.2
4x=0.8
/4
x=0.8/4
x=0.2= 20 cents = cost of apple
...
m.ananth@hotmail.ca